# Rep3D: Re-parameterize Large 3D Kernels with Low-Rank Receptive Modeling for Medical Imaging

### v1

@@ 5.2 Evaluation on Multi-Organ Segmentation
Beyond the ability to segment anatomical structures across scales, we furtehr evaluate Rep3D on the AMOS benchmark under the "train-from-scratch" setting for both CT and MRI modalities. On AMOS-CT, Rep3D achieves the best performance across all 15 evaluated anatomical structures, surpassing strong baselines including SwinUNETR, UNesT, and 3D UX-Net. Notably, Rep3D outperforms UNesT-B by 2.13% and RepOptimizer by 2.02% of average Dice score, while operating with fewer parameters than UNesT. On AMOS-MRI, a more challenging modality due to the variable range of contrast intensity and anatomical ambiguity, Rep3D maintains its superior performance, achieving an average Dice of 0.864, again outperforming all competing approaches. Compared to the best-performing transformer baseline (UNesT-B, 0.854) and convolutional baseline (3D UX-Net (k=21), 0.840), Rep3D delivers consistent improvements across nearly all organ classes, particularly in difficult regions such as the pancreas, gallbladder, and adrenal glands. These gains underscore the effectiveness of our spatially adaptive re-parameterization strategy in enhancing convergence and feature expressivity without increasing model complexity.
Table 3: Ablation Studies on the AMOS testing split
Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg Kernel=1×1×$\times 1\times 1$ 1 0.972 0.968 0.965 0.926 0.863 0.984 0.917 0.956 0.922 0.851 0.816 0.779 0.863 0.912 0.894 0.905 Kernel=3×3×$\times 3\times 3$ 3 0.970 0.966 0.960 0.930 0.863 0.984 0.935 0.958 0.924 0.859 0.827 0.758 0.862 0.908 0.892 0.906 Kernel=5×5×$\times 5\times 5$ 5 0.974 0.967 0.964 0.925 0.833 0.984 0.924 0.956 0.910 0.850 0.829 0.786 0.843 0.921 0.884 0.903 Kernel=7×7×$\times 7\times 7$ 7 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910 3D UX-Net (k=7) 0.966 0.959 0.951 0.903 0.833 0.980 0.910 0.950 0.913 0.830 0.805 0.756 0.846 0.897 0.863 0.890 3D UX-Net + LRBM 0.968 0.963 0.952 0.911 0.841 0.981 0.915 0.959 0.920 0.835 0.811 0.770 0.851 0.901 0.872 0.897
@@ 5.3 Ablation Studies
Effect of Network Depth for LRBM. To investigate the impact of architectural depth in the spatial modulation generator, we conduct an ablation study by varying the number of layers in the generator network used to produce the element-wise modulation mask in Rep3D (in supplementart material). Specifically, we compare shallow configurations (1-layer depthwise convolution) with deeper variants (2-layer and 3-layer depthwise convolution stacks), while keeping the total parameter count approximately constant. Our results show that the 2-layer design provides the best trade-off between representation flexibility and training stability. While the 1-layer generator lacks sufficient capacity to capture nuanced spatial priors, resulting in under-modulated gradient flow. The 3-layer version demonstrates a slight decrease of performance (from 0.910 to 0.899 Dice) and instability during training. This suggests that a lightweight, moderately deep generator is optimal for learning spatially adaptive convergence patterns without incurring additional complexity or over-parameterization.
Effect of Kernel Size in Spatial Bias Modeling. To further understand how kernel size affects segmentation performance across different anatomical structures, we analyzed organ-wise performance under varying kernel configurations: $1\times 1\times 1$, $3\times 3\times 3$, $5\times 5\times 5$, and $7\times 7\times 7$ used in Rep3D. All configurations share the same training protocol and re-parameterization setup, isolating the effect of kernel size alone. As shown in Table 4, the impact of kernel size varies across organs. While the $7\times 7\times 7$ kernel achieves the highest overall mean Dice score (0.910), smaller or boundary-sensitive organs (e.g., bladder, adrenal glands) benefit from small- or mid-size kernels such as $1\times 1\times 1$ or $5\times 5\times 5$. In contrast, large organs with strong spatial continuity (e.g., liver, spleen, aorta) show clear improvements with larger receptive fields. These results suggest that optimal kernel size is organ-dependent, influenced by factors such as spatial extent, anatomical context, and structural complexity. The superior performance of the $7\times 7\times 7$ variant reflects its ability to balance local detail and global context.
Effect of LRBM towards Other Network Architectures. To isolate the contribution of our proposed Low-Rank Bias Modeling (LRBM) module, we integrate it into a standard 3D UX-Net architecture (with fixed $7\times 7\times 7$ kernels) and compare its performance to the original baseline. As reported in Table 4, incorporating LRBM improves the average Dice score from 0.890 to 0.897, with consistent gains across multiple organs including pancreas, bladder, and adrenal glands. While the improvement may appear modest in aggregate, it is particularly noteworthy in anatomically challenging regions where gradient convergence is often unstable. For example, performance on the left adrenal gland increases from 0.756 to 0.770, and the duodenum improves from 0.846 to 0.851, suggesting that the learnable spatial bias improves optimization dynamics in fine-scale structures. These results confirm that our LRBM module offers a generalizable and plug-and-play mechanism for enhancing 3D segmentation backbones, even outside the full Rep3D framework.
6 Discussions & Limitations
@@
In this work, we introduced Rep3D, a re-parameterization framework that explicitly models spatial convergence dynamics in large kernel 3D convolutions. By linking effective receptive field (ERF) behavior with first-order optimization theory, we demonstrated that large convolution kernels naturally exhibit non-uniform learning dynamics, where central elements converge faster than peripheral ones. To address this, Rep3D integrates a learnable spatial prior via low-rank modulation, allowing the optimizer to differentially emphasize kernel regions with the distinctive characteristics of ERF during training. Our experiments across five diverse 3D segmentation benchmarks, confirm that Rep3D consistently improves performance over both transformer-based and convolution-based SOTA approaches, while maintaining a plain and efficient encoder design. The success of Rep3D reinforces several broader insights. First, spatially adaptive optimization is a promising direction for bridging inductive biases in CNNs with the dynamic learning capacity of attention-based models. Second, incorporating explicit ERF modeling into kernel design enables more efficient parameter usage, particularly in data-limited medical imaging scenarios. Moreover, our framework enhance network interpretability: the modulation masks can be visualized and aligned with ERF patterns, offering insights into how spatial understanding guides the learning of convolution kernels.
@@
While Rep3D demonstrates strong empirical performance across diverse 3D medical segmentation tasks, several limitations remain. First, although our learnable modulation mechanism introduces minimal architectural overhead, the training cost associated with large 3D kernels (e.g., $21\times 21\times 21$) remains nontrivial, particularly in memory-constrained GPU environments. Unlike 2D convolution kernels (i.e. MegEngine packages for 2D depthwise kernels), limited packages and approaches has been proposed to optimize the large kernel mechanism in 3D. This limits the batch size and input resolution during training, which can affect convergence and generalization. Future work could explore progressive training strategies, multi-resolution optimization, or low-resolution proxy supervision to alleviate this constraint while maintaining segmentation fidelity. Second, while our distance decay prior effectively guides spatial re-parameterization, its performance is inherently tied to the input volume resolution. In our experiments, we downsample 3D volumes to specific resolution (e.g., $1.5\times 1.5\times 2.0$ mm) to balance computation and efficency. However, we observe saturation effects when training at higher resolutions, where further improvements in image quality do not yield proportional gains in segmentation accuracy. This may be due to the spatial prior losing precision at finer scales. Adapting fine-grained spatial learnable prior could be another potential direction for future work.
@@ 7 Conclusion
Appendix A Data Preprocessing & Training Details
@@
Hyperparameters Direct Training Encoder Stage 4 Layer-wise Channel $48,96,192,384$ Hidden Dimensions $768$ Patch Size $96\times 96\times 96$ No. of Sub-volumes Cropped 2 Training Steps 60000 Batch Size 2 AdamW $\epsilon$ $81e-$ 8 AdamW $\beta$ ($0.9,0.999$) Peak Learning Rate $41e-$ 4 Learning Rate Scheduler ReduceLROnPlateau Factor & Patience 0.9, 10 Dropout X Weight Decay 0.08 Data Augmentation Intensity Shift, Rotation, Scaling Cropped Foreground Intensity Offset 0.1 Rotation Degree $-30^{\circ}$ to $+30^{\circ}$ Scaling Factor x: 0.1, y: 0.1, z: 0.1
We apply hierarchical steps for data preprocessing: 1) intensity clipping is applied to further enhance the contrast of soft tissue (AMOS CT, KiTS, MSD Pancreas:{min:-175, max:250}; MSD Hepatic Vessel:{min:0, max:230}); AMOS MRI:{min:0, max:1000}. 2) Intensity normalization is performed after clipping for each volume and use min-max normalization: $(X-X_{1})/(X_{99}-X_{1})$ to normalize the intensity value between 0 and 1, where $X_{p}$ denote as the $p_{th}$ percentile of intensity in $X$. We then perform downsampling to certain voxel spacing (i.e. AMOS CT, MSD hepatic vessels, MSD Pancreas and KiTS: $1.5\times 1.5\times 2.0$, AMOS MRI: $1.0\times 1.0\times 1.0$) randomly crop sub-volumes with size $96\times 96\times 96$ at the foreground and perform data augmentations, including rotations, intensity shifting, and scaling (scaling factor: 0.1). All training processes with Rep3D are optimized with either Stochastic Gradient Descent (SGD) or AdamW optimizer. We trained all models for 60000 steps using a learning rate of $0.0001$ on an NVIDIA A100 GPU across all datasets. One epoch takes approximately about 9 minute for KiTS, 5 minutes for MSD Pancreas, 12 minutes for MSD hepatic vessels, 7 minutes for AMOS CT and 1 minute for AMOS MRI, respectively. We further summarize all the training parameters with Table 1 in Supplementary Material.
Appendix B Datasets Details

### v2

@@ 5.2 Evaluation on Multi-Organ Segmentation
Beyond the ability to segment anatomical structures across scales, we further evaluate Rep3D on the AMOS benchmark under the “train-from-scratch” setting for both CT and MRI modalities. On AMOS-CT, Rep3D achieves the best performance across all 15 evaluated anatomical structures, surpassing strong baselines including SwinUNETR, UNesT, and 3D UX-Net. Notably, Rep3D outperforms UNesT-B by 2.13% and RepOptimizer by 2.02% of average Dice score, while operating with fewer parameters than UNesT. On AMOS-MRI, a more challenging modality due to the variable range of contrast intensity and anatomical ambiguity, Rep3D maintains its superior performance, achieving an average Dice of 0.864, again outperforming all competing approaches. Compared to the best-performing transformer baseline (UNesT-B, 0.854) and convolutional baseline (3D UX-Net (k=21), 0.840), Rep3D delivers consistent improvements across nearly all organ classes, particularly in difficult regions such as the pancreas, gallbladder, and adrenal glands. These gains underscore the effectiveness of our spatially adaptive re-parameterization strategy in enhancing convergence and feature expressivity without increasing model complexity.
@@ 5.3 Ablation Studies
Isolation of Spatial Modulation (LRBM vs. Vanilla). To strictly isolate the contribution of our proposed Low-Rank Receptive Bias Modeling (LRBM), we compare Rep3D against a “Vanilla Rep” baseline (Appendix Table 7). This baseline utilizes the identical plain encoder architecture (a single $21\times 21\times 21$ branch per block) but is trained using standard stochastic gradient descent without spatial modulation. The results are decisive: the Vanilla Rep baseline exhibits significant optimization instability, characterized by high variance in training loss and consistently lower Dice scores. This empirical failure confirms that the “plain” large-kernel architecture is inherently unstable on volumetric data; our learnable spatial prior is not merely an enhancement, but a necessary condition for convergence in this regime.
Optimal Capacity for Bias Modeling (Generator Depth). We investigate the representational capacity required to model the spatial gradient prior effectively. By varying the depth of the generator network $f_{\theta}$ (from 1 to 3 layers), we observe a clear trade-off between expressivity and optimizability. The 1-layer generator proves insufficient, likely because a linear mapping cannot capture the complex, non-isotropic diffusion patterns of the effective receptive field. Conversely, the 3-layer variant introduces optimization difficulties, leading to a slight performance degradation ($0.910\to 0.899$ Dice). Our results identify the 2-layer configuration as the optimal design, providing sufficient non-linearity to model nuanced spatial priors while maintaining a lightweight footprint that does not impede gradient flow.
Local vs. Global Context in Bias Generation (Generator Kernel Size). While the Rep3D backbone utilizes a massive $21\times 21\times 21$ kernel to capture global semantic context, the spatial bias generator may operate on a different scale. We analyzed the generator’s kernel size ($K_{G}\in\{1^{3},3^{3},5^{3},7^{3}\}$) to determine the optimal receptive field for determining gradient weights. As shown in Table 3, the impact is organ-dependent. Small generator kernels ($1^{3}$) favor boundary-sensitive organs (e.g., bladder), suggesting that fine-grained local statistics are crucial for edge delineation. However, the $7\times 7\times 7$ generator achieves the highest overall Dice (0.910), indicating that a larger receptive field is the best for estimating the spatial prior. This suggests that the decision to up-weight or down-weight a central pixel should be informed by its immediate neighborhood, even if the backbone filter itself spans a much larger volume.
Universality of LRBM (Integration with 3D UX-Net). Finally, to demonstrate that our findings are not specific to the Rep3D architecture, we integrate the LRBM module into a standard 3D UX-Net (which uses fixed $7\times 7\times 7$ kernels). As reported in Table 3, this integration improves the average Dice score from 0.890 to 0.897. The gains are particularly pronounced in anatomically challenging regions with irregular shapes, such as the left adrenal gland ($0.756\to 0.770$) and the duodenum ($0.846\to 0.851$). These improvements confirm that spatially adaptive learning rates provide a robust inductive bias that aids convergence in hard-to-segment regions, functioning as a plug-and-play enhancement for existing volumetric segmentation backbones.
6 Discussion & Limitations
@@
In this work, we introduced Rep3D, a re-parameterization framework that explicitly models spatial convergence dynamics in large kernel 3D convolutions. By linking effective receptive field (ERF) behavior with first-order optimization theory, we demonstrated that large convolution kernels naturally exhibit non-uniform learning dynamics, where central elements converge faster than peripheral ones. To address this, Rep3D integrates a learnable spatial prior via low-rank modulation, allowing the optimizer to differentially emphasize kernel regions with the distinctive characteristics of ERF during training. Our experiments across five diverse 3D segmentation benchmarks, confirm that Rep3D consistently improves performance over both transformer-based and convolution-based SOTA approaches, while maintaining a plain and efficient encoder design. The success of Rep3D reinforces several broader insights. First, spatially adaptive optimization is a promising direction for bridging inductive biases in CNNs with the dynamic learning capacity of attention-based models. Second, incorporating explicit ERF modeling into kernel design enables more efficient parameter usage, particularly in data-limited medical imaging scenarios. Moreover, our framework enhance network interpretability: the modulation masks can be visualized and aligned with ERF pattern as demonstrated in Figure 3, offering insights into how spatial understanding guides the learning of convolution kernels.
@@
While Rep3D demonstrates strong empirical performance across diverse 3D medical segmentation tasks, several limitations remain. First, although our learnable modulation mechanism introduces minimal architectural overhead, the training cost associated with large 3D kernels (e.g., $21\times 21\times 21$) remains nontrivial, particularly in memory-constrained GPU environments. Unlike 2D convolution kernels (i.e. MegEngine packages for 2D depthwise kernels), limited packages and approaches have been proposed to optimize the large kernel mechanism in 3D. This limits the batch size and input resolution during training, which can affect convergence and generalization. Future work could explore progressive training strategies, multi-resolution optimization, or low-resolution proxy supervision to alleviate this constraint while maintaining segmentation fidelity. Second, while our distance decay prior effectively guides spatial re-parameterization, its performance is inherently tied to the input volume resolution. In our experiments, we downsample 3D volumes to specific resolution (e.g., $1.5\times 1.5\times 2.0$ mm) to balance computation and efficiency. However, we observe saturation effects when training at higher resolutions, where further improvements in image quality do not yield proportional gains in segmentation accuracy. This may be due to the spatial prior losing precision at finer scales. Adapting fine-grained spatial learnable prior could be another potential direction for future work.
@@ 7 Conclusion
Appendix A Rep3D Motivations from Experiments: Naively Scaling 3D Convolution Kernels
Recent advances in medical image segmentation have highlighted the importance of large receptive fields for capturing long-range spatial dependencies in volumetric data. Motivated by this, there has been a growing trend toward enlarging convolutional kernel sizes in 3D CNN architectures, such as 3D UX-Net (Lee et al., 2022) and RepUX-Net (Lee et al., 2023), which attempt to mimic the global context modeling capabilities of transformers while retaining the inductive biases of CNNs.
However, the straightforward enlargement of kernel size introduces several practical and theoretical challenges:
  • Optimization Instability: Large convolutional kernels suffer from slow or uneven convergence, particularly in the outer kernel regions, which are rarely activated in early training. This leads to ineffective utilization of capacity and suboptimal learning behavior.
  • Degrading Performance: Empirically, simply increasing the kernel size does not guarantee improved performance. Beyond a certain scale, performance tends to saturate or even degrade.
  • Inefficient Parameter Usage: Naïve kernel scaling introduces a quadratic (especially in 3D) growth in the number of parameters, making training inefficient and difficult to regularize.
To illustrate these effects empirically, we conduct a systematic ablation study on the AMOS dataset using the 3D UX-Net encoder as a backbone. 3D UX-Net is an ideal starting point because it already leverages larger kernels ($7\times 7\times 7$) in its baseline. We vary the convolutional kernel size from $3\times 3\times 3$ to $21\times 21\times 21$, keeping all other architectural and training settings fixed. The results, shown in Table 4 , confirm our hypothesis.
Table 4: Impact of kernel size on segmentation performance using 3D UX-Net encoder on the AMOS dataset.
Kernel Size Avg. Dice Score
$3\times 3\times 3$ 0.881
$5\times 5\times 5$ 0.885
$7\times 7\times 7$ 0.890
$9\times 9\times 9$ 0.891
$11\times 11\times 11$ 0.893
$13\times 13\times 13$ 0.894
$15\times 15\times 15$ 0.895
$17\times 17\times 17$ 0.893
$19\times 19\times 19$ 0.893
$21\times 21\times 21$ 0.891
As seen in Table 4, performance initially improves with increasing kernel size, peaking at $15\times 15\times 15$. However, further enlargements yield diminishing or even negative returns, despite increased computational cost. These findings reveal a key limitation of naïve kernel enlargement: Although it increases theoretical receptive field, it fails to translate into meaningful gains due to the lack of spatially adaptive optimization.
A.1 Theoretical extension from SGD to Adam/AdamW
We provide a rigorous derivation of how the Constant-Scale Linear Addition (CSLA) block induces a spatially varying learning rate, specifically under adaptive optimization algorithms like Adam (Kingma & Ba, 2014).
1. The Normalization Cancellation Effect Consider the update rule for a single parameter $w$ in Adam. The update $\Delta w_{t}$ at step $t$ is given by:
$\Delta w_{t}=-\eta\cdot\frac{m_{t}}{\sqrt{v_{t}}+\epsilon},$ (19)
where $m_{t}$ is the first moment (momentum) and $v_{t}$ is the second moment (variance) of the gradients. In a re-parameterized branch scaled by $\alpha$ (i.e., $w_{\text{branch}}=\alpha w$), the gradient is scaled by $\alpha$: $g\leftarrow\alpha g$. Consequently, the moments scale as $m_{t}\leftarrow\alpha m_{t}$ and $v_{t}\leftarrow\alpha^{2}v_{t}$. Substituting these into the update rule:
$\Delta w_{\text{branch}}=-\eta\cdot\frac{\alpha m_{t}}{\sqrt{\alpha^{2}v_{t}}+\epsilon}\approx-\eta\cdot\frac{m_{t}}{\sqrt{v_{t}}}.$ (20)
As noted, the scaling factor $\alpha$ cancels out. This implies that the magnitude of the update step for any single branch is largely independent of its scaling factor $\alpha$. It is effectively governed only by the global learning rate $\eta$.
2. Structural Superposition and Effective Step Size However, such derivation lies in analyzing branches in isolation. The effective kernel $W^{\prime}$ used during inference is the structural superposition of the branches:
$W^{\prime}=\alpha_{L}W_{L}+\alpha_{S}\mathcal{P}(W_{S}),$ (21)
where $\mathcal{P}(\cdot)$ represents the zero-padding operator that aligns the small kernel to the spatial dimensions of the large kernel.
The effective update $\Delta W^{\prime}$ is the linear combination of the branch updates:
$\Delta W^{\prime}(x)=\alpha_{L}\Delta W_{L}(x)+\alpha_{S}\Delta W_{S}(x).$ (22)
We can further analyze the magnitude of this effective update across spatial positions $x$:
Case A: Peripheral Region ($x\in\text{Periphery}$) These spatial positions exist only in the support of $W_{L}$. The effective update is:
$\Delta W^{\prime}_{\text{peri}}=\alpha_{L}\Delta W_{L}\approx\alpha_{L}\cdot\eta\cdot\vec{u}_{L},$ (23)
where $\vec{u}_{L}$ is the normalized update direction. The step size magnitude is proportional to one unit of update.
Case B: Central Region ($x\in\text{Center}$) These spatial positions exist in the support of both $W_{L}$ and $W_{S}$. The effective update is the superposition of both branches:
$\Delta W^{\prime}_{\text{center}}=\alpha_{L}\Delta W_{L}+\alpha_{S}\Delta W_{S}\approx\alpha_{L}(\eta\vec{u}_{L})+\alpha_{S}(\eta\vec{u}_{S}).$ (24)
Assuming the gradient directions $\vec{u}_{L}$ and $\vec{u}_{S}$ are not perfectly orthogonal (they are optimizing the same loss on the same features), these vectors constructively interfere. Crucially, even if Adam normalizes $\Delta W_{L}$ and $\Delta W_{S}$ to have similar magnitudes, the center receives the sum of two update vectors, while the periphery receives only one.
3. The Induced Spatial Prior This derivation proves that structurally re-parameterized blocks inherently induce a spatial learning rate field $\lambda(x)$:
$\lVert\Delta W^{\prime}(x)\rVert\propto\begin{cases}C_{1}\cdot\eta&\text{if }x\in\text{Periphery}\\ (C_{1}+C_{2})\cdot\eta&\text{if }x\in\text{Center}\end{cases}$ (25)
where $C_{1},C_{2}>0$ are constants derived from the structural combination. This confirms that optimization is faster at the center and slower at the periphery, aligning with the ”Gaussian” diffusion pattern of Effective Receptive Fields (ERFs), demonstrating similar outcome with the SGD optimizer.
A.2 Data Preprocessing & Training Details
@@
Hyperparameters Direct Training Encoder Stage 4 Layer-wise Channel $48,96,192,384$ Hidden Dimensions $768$ Patch Size $96\times 96\times 96$ No. of Sub-volumes Cropped 2 Training Steps 60000 Batch Size 1 AdamW $\epsilon$ 1e81e-8 AdamW $\beta$ ($0.9,0.999$) Peak Learning Rate 1e41e-4 Learning Rate Scheduler ReduceLROnPlateau Factor & Patience 0.9, 10 Dropout X Weight Decay 0.08 Data Augmentation Intensity Shift, Rotation, Scaling Cropped Foreground Intensity Offset 0.1 Rotation Degree $30^{\circ}$ to $30^{\circ}$ Scaling Factor x: 0.1, y: 0.1, z: 0.1
We apply hierarchical steps for data preprocessing: 1) intensity clipping is applied to further enhance the contrast of soft tissue (AMOS CT, KiTS, MSD Pancreas:{min:-175, max:250}; MSD Hepatic Vessel:{min:0, max:230}); AMOS MRI:{min:0, max:1000}. 2) Intensity normalization is performed after clipping for each volume and use min-max normalization: $(X-X_{1})/(X_{99}-X_{1})$ to normalize the intensity value between 0 and 1, where $X_{p}$ denote as the $p_{th}$ percentile of intensity in $X$. We then perform downsampling to certain voxel spacing (i.e. AMOS CT, MSD hepatic vessels, MSD Pancreas and KiTS: $1.5\times 1.5\times 2.0$, AMOS MRI: $1.0\times 1.0\times 1.0$) randomly crop sub-volumes with size $96\times 96\times 96$ at the foreground and perform data augmentations, including rotations, intensity shifting, and scaling (scaling factor: 0.1). All training processes with Rep3D are optimized with either Stochastic Gradient Descent (SGD) or AdamW optimizer. We trained all models for 60000 steps using a learning rate of $0.0001$ on an NVIDIA A100 GPU across all datasets. One epoch takes approximately about 9 minutes for KiTS, 5 minutes for MSD Pancreas, 12 minutes for MSD hepatic vessels, 7 minutes for AMOS CT and 1 minute for AMOS MRI, respectively.
All experiments are conducted under a direct supervised learning setting. For the KiTS and MSD datasets, we employ a 5-fold cross-validation strategy using an 80%/10%/10% split for training, validation, and testing, respectively. For the AMOS dataset, we use a fixed single split with the same partitioning ratio. Details on training procedures and preprocessing protocols are provided in the supplementary material. Our proposed re-parameterization approach Rep3D, is benchmarked against both convolutional and transformer-based state-of-the-art (SOTA) methods for 3D medical image segmentation. For nnUNet (Isensee et al., 2021), we evaluate performance across two different training schedules to account for fairness, since Rep3D is trained with 60,000 iterations (approximately equivalent to 350 epochs). Therefore, we have provided performance with partial scheduled (350 epochs) and full scheduled (1000 epochs) to demonstrate our model generalizability.
A.3 Datasets Details
We have leverage four challenging public datasets across different scales: 1) AMOS22 (MICCAI 2022 Abdominal Multi-organ Segmentation Challenge) (Ji et al., 2022): Comprises 200 multi-contrast abdominal CT scans with 15 organ-level anatomical labels and 33 MRI scans with 13 organ-level anatomical labels for comprehensive abdominal segmentation, 2) KiTS21 (MICCAI 2021 Kidney Tumor Segmentation Challenge) (Heller et al., 2019): Includes 210 contrast-enhanced abdominal CT scans from the University of Minnesota Medical Center (2010–2018), with manual annotations for kidney, tumor, and cyst, 3) MSD Pancreas (Medical Segmentation Decathlon) (Antonelli et al., 2022): Contains 282 abdominal contrast-enhanced CT scans annotated for both pancreas and pancreatic tumor segmentation, and 4) MSD Hepatic Vessel (Medical Segmentation Decathlon) (Antonelli et al., 2022): Contains 303 abdominal CT scans annotated for hepatic vessel and associated tumor segmentation.